# Proving the Limited Scalability of Centralized Distributed Optimization via a New Lower Bound Construction

### v1

@@ Abstract
We consider centralized distributed optimization in the classical federated learning setup, where $n$ workers jointly find an $\varepsilon$–stationary point of an $L$–smooth, $d$–dimensional nonconvex function $f$, having access only to unbiased stochastic gradients with variance $\sigma^{2}$. Each worker requires at most $h$ seconds to compute a stochastic gradient, and the communication times from the server to the workers and from the workers to the server are $\tau_{\textnormal{s}}$ and $\tau_{\textnormal{w}}$ seconds per coordinate, respectively. One of the main motivations for distributed optimization is to achieve scalability with respect to $n$. For instance, it is well known that the distributed version of SGD has a variance-dependent runtime term $\nicefrac{{h\sigma^{2}L\Delta}}{{n\varepsilon^{2}}},$ which improves with the number of workers $n,$ where $\Delta:=f(x^{0})-f^{*},$ and $x^{0}\in\mathbb{R}^{d}$ is the starting point. Similarly, using unbiased sparsification compressors, it is possible to reduce both the variance-dependent runtime term and the communication runtime term from $\tau_{\textnormal{w}}d\nicefrac{{L\Delta}}{{\varepsilon}}$ to $\nicefrac{{\tau_{\textnormal{w}}dL\Delta}}{{n\varepsilon}}+\sqrt{\nicefrac{{\tau_{\textnormal{w}}dh\sigma^{2}}}{{n\varepsilon}}}\cdot\nicefrac{{L\Delta}}{{\varepsilon}},$ which also benefits from increasing $n.$ However, once we account for the communication from the server to the workers $\tau_{\textnormal{s}}$, we prove that it becomes infeasible to design a method using unbiased random sparsification compressors that scales both the server-side communication runtime term $\tau_{\textnormal{s}}d\nicefrac{{L\Delta}}{{\varepsilon}}$ and the variance-dependent runtime term $\nicefrac{{h\sigma^{2}L\Delta}}{{\varepsilon^{2}}},$ better than poly-logarithmically in $n$, even in the homogeneous (i.i.d.) case, where all workers access the same function or distribution. Indeed, when $\tau_{\textnormal{s}}\simeq\tau_{\textnormal{w}},$ our lower bound is $\tilde{\Omega}\left(\min\left\{h\left(\frac{\sigma^{2}}{n\varepsilon}+1\right)\frac{L\Delta}{\varepsilon}+{\color[rgb]{0.92578125,0.578125,0.0546875}\definecolor[named]{pgfstrokecolor}{rgb}{0.92578125,0.578125,0.0546875}\tau_{\textnormal{s}}d\frac{L\Delta}{\varepsilon}},\;h\frac{L\Delta}{\varepsilon}+{\color[rgb]{0.92578125,0.578125,0.0546875}\definecolor[named]{pgfstrokecolor}{rgb}{0.92578125,0.578125,0.0546875}h\frac{\sigma^{2}L\Delta}{\varepsilon^{2}}}\right\}\right).$ To establish this result, we construct a new “worst-case” function and develop a new lower bound framework that reduces the analysis to the concentration of a random sum, for which we prove a concentration bound. These results reveal fundamental limitations in scaling distributed optimization, even under the homogeneous (i.i.d.) assumption.
@@ 1 Introduction
$f$ is differentiable & $L$–smooth, i.e., $\left\|\nabla f(x)-\nabla f(y)\right\|\leq L\left\|x-y\right\|$, $\forall x,y\in\mathbb{R}^{d}.$
@@
The goal in the nonconvex world is to find an $\varepsilon$–stationary point, a (random) point $\bar{x}\in\mathbb{R}^{d}$ such that ${\rm\mathbb{E}}[\|\nabla f(\bar{x})\|^{2}]\leq\varepsilon$ (Nemirovskij & Yudin, 1983; Murty & Kabadi, 1985). We also consider a realistic computation and communication scenario:
@@ 1.1 Related work
1. Communication is free. Let us temporarily assume that communication does not take time, i.e., $\tau_{\textnormal{s}}=0$ and $\tau_{\textnormal{w}}=0.$ Then, in this scenario, the theoretically fastest strategy is to run the Synchronized SGD method, i.e., $x^{k+1}=x^{k}-\frac{\gamma}{n}\sum_{i=1}^{n}\nabla f(x^{k};\xi^{k}_{i}),$ where $\gamma=\Theta(\min\{\nicefrac{{1}}{{L}},\nicefrac{{\varepsilon n}}{{L\sigma^{2}}}\}),$$\{\xi^{k}_{i}\}$ are i.i.d., and $\{\nabla f(x^{k};\xi^{k}_{i})\}$ are computed in parallel by the workers, which send to the server that aggregates and calculates $x^{k+1}.$ One can show that the time complexity of this method is $\mathcal{O}\bigl{(}h\bigl{(}\frac{L\Delta}{\varepsilon}+\frac{\sigma^{2}L\Delta}{n\varepsilon^{2}}\bigr{)}\bigr{)},$ because it requires $\mathcal{O}\bigl{(}\frac{L\Delta}{\varepsilon}+\frac{\sigma^{2}L\Delta}{n\varepsilon^{2}}\bigr{)}$ iterations (Lan, 2020), and each iteration takes at most $h$ seconds due to Assumption 1.4. Moreover, this result is optimal and can not be improved (Arjevani et al., 2022; Tyurin & Richtárik, 2023b).
@@
$x^{k+1}=x^{k}-\frac{\gamma}{n}\sum\limits_{i=1}^{n}\frac{1}{b}\sum\limits_{j=1}^{b}\nabla f(x^{k};\xi^{k}_{ij})$ (Batch Synchronized SGD)
@@
$x^{k+1}=x^{k}-\frac{\gamma}{nbm}\sum\limits_{i=1}^{n}\sum\limits_{k=1}^{m}\mathcal{C}_{ik}\left(\sum\limits_{j=1}^{b}\nabla f(x^{k};\xi^{k}_{ij})\right),$ (Batch QSGD)
@@
A high-level explanation for why the dependence on $d$ improves with $n$ is that all workers use i.i.d. and unbiased compressors $\{\mathcal{C}_{ik}\},$ which allow them to collaboratively explore more coordinates. This effect is similar to Synchronized SGD, where the variance ${\rm\mathbb{E}}_{\xi}[\|\frac{1}{n}\sum_{i=1}^{n}\nabla f(x;\xi^{k}_{i})-\nabla f(x)\|^{2}]\leq\frac{\sigma^{2}}{n}$ also improves with $n.$ There are many other compressed methods that also improve with $n,$ including DIANA (Mishchenko et al., 2019), Accelerated DIANA (Li et al., 2020), MARINA (Gorbunov et al., 2021), DASHA (Tyurin & Richtárik, 2023a), and FRECON (Zhao et al., 2021).
@@ 1.2 Contributions
Theorem 1.6 (Informal Formulation of Theorems 4.2 and D.1).
@@
Because of the $\min$, the bound shows that it is possible to improve either the dependence on $d$ or the dependence on $\nicefrac{{\sigma^{2}}}{{\varepsilon}}$ as the number of workers $n$ increases, but not both simultaneously. The lower bound is matched either by Batch QSGD or by the non-distributed SGD method (without any communication or cooperation). Moreover, if $\tau_{\textnormal{s}}\simeq\tau_{\textnormal{w}},$ the lower bound becomes $\tilde{\Omega}\left(\min\left\{h\left(\frac{\sigma^{2}}{n\varepsilon}+1\right)\frac{L\Delta}{\varepsilon}+{\color[rgb]{0.92578125,0.578125,0.0546875}\definecolor[named]{pgfstrokecolor}{rgb}{0.92578125,0.578125,0.0546875}\tau_{\textnormal{s}}d\frac{L\Delta}{\varepsilon}},h\frac{L\Delta}{\varepsilon}+{\color[rgb]{0.92578125,0.578125,0.0546875}\definecolor[named]{pgfstrokecolor}{rgb}{0.92578125,0.578125,0.0546875}h\frac{\sigma^{2}L\Delta}{\varepsilon^{2}}}\right\}\right),$ which can be matched by Batch Synchronized SGD with the complexity (2) (without compression) or by the non-distributed SGD method.
@@
Improved analysis when $\tau_{\textnormal{w}}>0.$ To obtain the complete lower bound, we extended and improved the result by Tyurin et al. (2024), which was limited for our scenario and required additional modifications to finally obtain (5) (see Sections D and 5 for details).
@@ 2 Preliminaries
For better comprehension of our new idea, we now present arguably one of the most important worst-case functions by Carmon et al. (2020), which is widely used to prove lower bounds in nonconvex optimization. It has been used by Arjevani et al. (2022; 2020) to derive lower bounds in the stochastic setting, by Lu & De Sa (2021) in the decentralized setting, by Tyurin & Richtárik (2023b; 2024); Tyurin et al. (2024) in the asynchronous setting, by Huang et al. (2022) to show the lower iteration bound for unidirectional compressed methods, and by Li et al. (2021) in problems with a nonconvex-strongly-concave structure.
@@
[…]
Notice that this function has a “chain–like” structure. If a method starts from $x^{0}=0$ and computes the gradient of $F_{T},$ then the gradient will have a non-zero value only in the first coordinate (use that $\Psi(0)=\Psi^{\prime}(0)=0$). Thus, by computing a single gradient, any “reasonable” method can “discover” at most one coordinate. At the same time, if the method wants to find an $\varepsilon$–stationary point, it should eventually discover the $T$th coordinate. These two facts imply that every “reasonable” method should compute the gradient of $F_{T}$ at least $T$ times. In the construction, Carmon et al. (2020) take $T=\Theta\left(\frac{L\Delta}{\varepsilon}\right).$ This construction is a “more technical” version of the celebrated quartic optimization construction from (Nesterov, 2018), which has similar properties. Let us define $\textnormal{prog}(x):=\max\{i~{}\geq~{}0\,|\,x_{i}\neq 0\}\,\,(x_{0}\equiv 1),$ then the following lemma is a formalization of the described properties.
@@ 2.1 Family of distributed methods
In our lower bound, we focus on the family of methods described by Protocol 1. This protocol takes an algorithm as input and runs the standard functions of the workers and the server: the workers compute stochastic gradients locally, the server aggregates them asynchronously and in parallel, and sends compressed information back based on the local information. For now, notice that we ignore the communication times from the workers to the server in Protocol 1, assuming that $\tau_{\textnormal{w}}=0.$ If we start considering this cost, then the following lower bounds can only increase.
@@
This is the standard assumption (Carmon et al., 2020) that generalizes the family of methods working with the span of vectors (Nesterov, 2018) and holds for the majority of methods, including GD, Adam (Kingma & Ba, 2015), DORE (Liu et al., 2020), EF21-P (Gruntkowska et al., 2023), MARINA-P, and M3 (Gruntkowska et al., 2024).
@@ 2.2 Previous Lower Bound in the Heterogeneous Setting
4:  Run the following two loops in parallel. The first on the workers. The second on the server.
@@
6:     while True do
7:        Algorithm $A$ calculates a new point $x$ based on local information $I_{i}$ (takes 00 seconds)
8:        Calculate a stochastic gradient [^3] In the heterogeneous setup (Section 2.2), worker $i$ computes $\nabla f_{i}(x;\xi),$ where $f_{i}$ is its local function.$\nabla f(x;\xi),$ $\quad\xi\sim\mathcal{D}_{\xi}$ (takes $h$ seconds)
9:        Add $\nabla f(x;\xi)$ to $I_{i}$ (takes 00 seconds)
10:        Share $\nabla f(x;\xi)$ with the server: add $\nabla f(x;\xi)$ to $I$ (takes 00 seconds)
11:     end while
12:  end for
13:  for $i=1,\dots,n$ (in parallel on the server) do
14:     for $k=0,1,\dots$ do
15:        Algorithm $A$ calculates new points $s_{i}^{k}$ based on local information $I$ (takes 00 seconds)
16:        Algorithm $A$ compresses the points $\mathcal{C}^{k}_{i}(s_{i}^{k})$ (takes 00 seconds)
17:        Send $\mathcal{C}^{k}_{i}(s_{i}^{k})$ to the worker $i$: add to $\mathcal{C}^{k}_{i}(s_{i}^{k})$ to $I_{i}$ (takes $\tau_{\textnormal{s}}\times P^{k}_{i}$ seconds, where $P^{k}_{i}$ is the number of coordinates retained by $\mathcal{C}^{k}_{i}(s_{i}^{k})$)
@@
19:  end for
@@
[…]

### v2

@@ Abstract
We consider centralized distributed optimization in the classical federated learning setup, where $n$ workers jointly find an $\varepsilon$–stationary point of an LL–smooth, dd–dimensional nonconvex function $f$, having access only to unbiased stochastic gradients with variance $\sigma^{2}$. Each worker requires at most $h$ seconds to compute a stochastic gradient, and the communication times from the server to the workers and from the workers to the server are $\tau_{\textnormal{s}}$ and $\tau_{\textnormal{w}}$ seconds per coordinate, respectively. One of the main motivations for distributed optimization is to achieve scalability with respect to $n$. For instance, it is well known that the distributed version of SGD has a variance-dependent runtime term $\nicefrac{{h\sigma^{2}L\Delta}}{{n\varepsilon^{2}}},$ which improves with the number of workers $n,$ where $\Delta:=f(x^{0})-f^{*},$ and $x^{0}\in\mathbb{R}^{d}$ is the starting point. Similarly, using unbiased sparsification compressors, it is possible to reduce both the variance-dependent runtime term and the communication runtime term from $\tau_{\textnormal{w}}d\nicefrac{{L\Delta}}{{\varepsilon}}$ to $\nicefrac{{\tau_{\textnormal{w}}dL\Delta}}{{n\varepsilon}}+\sqrt{\nicefrac{{\tau_{\textnormal{w}}dh\sigma^{2}}}{{n\varepsilon}}}\cdot\nicefrac{{L\Delta}}{{\varepsilon}},$ which also benefits from increasing $n.$ However, once we account for the communication from the server to the workers $\tau_{\textnormal{s}}$, we prove that it becomes infeasible to design a method using unbiased random sparsification compressors that scales both the server-side communication runtime term $\tau_{\textnormal{s}}d\nicefrac{{L\Delta}}{{\varepsilon}}$ and the variance-dependent runtime term $\nicefrac{{h\sigma^{2}L\Delta}}{{\varepsilon^{2}}},$ better than poly-logarithmically in $n$, even in the homogeneous (i.i.d.) case, where all workers access the same function or distribution. Indeed, when $\tau_{\textnormal{s}}\simeq\tau_{\textnormal{w}},$ our lower bound is $\tilde{\Omega}\left(\min\left\{h\left(\frac{\sigma^{2}}{n\varepsilon}+1\right)\frac{L\Delta}{\varepsilon}+{\color[rgb]{0.92578125,0.578125,0.0546875}\definecolor[named]{pgfstrokecolor}{rgb}{0.92578125,0.578125,0.0546875}\tau_{\textnormal{s}}d\frac{L\Delta}{\varepsilon}},\;h\frac{L\Delta}{\varepsilon}+{\color[rgb]{0.92578125,0.578125,0.0546875}\definecolor[named]{pgfstrokecolor}{rgb}{0.92578125,0.578125,0.0546875}h\frac{\sigma^{2}L\Delta}{\varepsilon^{2}}}\right\}\right).$ To establish this result, we construct a new “worst-case” function and develop a new lower bound framework that reduces the analysis to the concentration of a random sum, for which we prove a concentration bound. These results reveal fundamental limitations in scaling distributed optimization, even under the homogeneous (i.i.d.) assumption.
@@ 1 Introduction
$f$ is differentiable & LL–smooth, i.e., $\left\|\nabla f(x)-\nabla f(y)\right\|\leq L\left\|x-y\right\|$, $\forall x,y\in\mathbb{R}^{d}.$
@@
The goal in the nonconvex world is to find an $\varepsilon$–stationary point, a (random) point $\bar{x}\in\mathbb{R}^{d}$ such that ${\rm\mathbb{E}}[\|\nabla f(\bar{x})\|^{2}]\leq\varepsilon$ (Nemirovskij and Yudin, 1983; Murty and Kabadi, 1985). We also consider a realistic computation and communication scenario:
@@ 1.1 Related work
1. Communication is free. Let us temporarily assume that communication does not take time, i.e., $\tau_{\textnormal{s}}=0$ and $\tau_{\textnormal{w}}=0.$ Then, in this scenario, the theoretically fastest strategy is to run the Synchronized SGD method, i.e., $x^{k+1}=x^{k}-\frac{\gamma}{n}\sum_{i=1}^{n}\nabla f(x^{k};\xi^{k}_{i}),$ where $\gamma=\Theta(\min\{\nicefrac{{1}}{{L}},\nicefrac{{\varepsilon n}}{{L\sigma^{2}}}\}),$ $\{\xi^{k}_{i}\}$ are i.i.d., and $\{\nabla f(x^{k};\xi^{k}_{i})\}$ are computed in parallel by the workers, which send to the server that aggregates and calculates $x^{k+1}.$ One can show that the time complexity of this method is $\mathcal{O}\bigl(h\bigl(\frac{L\Delta}{\varepsilon}+\frac{\sigma^{2}L\Delta}{n\varepsilon^{2}}\bigr)\bigr),$ because it requires $\mathcal{O}\bigl(\frac{L\Delta}{\varepsilon}+\frac{\sigma^{2}L\Delta}{n\varepsilon^{2}}\bigr)$ iterations (Lan, 2020), and each iteration takes at most $h$ seconds due to Assumption 1.4. Moreover, this result is optimal and can not be improved (Arjevani et al., 2022; Tyurin and Richtárik, 2023b).
@@
[…]
@@
A standard example of an unbiased compressor is RandKK $\in\mathbb{U}(\nicefrac{{d}}{{K}}-1),$ which selects $K$ random coordinates of the input vector $x$, scales them by $\nicefrac{{d}}{{K}}$, and sets the remaining coordinates to zero (see Def. C.1). Numerous other examples of unbiased compressors have been explored in the literature (Beznosikov et al., 2020; Xu et al., 2021; Horváth et al., 2022; Szlendak et al., 2021). Using the seminal ideas (Seide et al., 2014), we can construct a modified version of QSGD (Alistarh et al., 2017) (special case of Shadowheart SGD from (Tyurin et al., 2024)), which we call Batch QSGD:
@@
A high-level explanation for why the dependence on $d$ improves with $n$ is that all workers use i.i.d. and unbiased compressors $\{\mathcal{C}_{ik}\},$ which allow them to collaboratively explore more coordinates. This effect is similar to Synchronized SGD, where the variance ${\rm\mathbb{E}}_{\xi}[\|\frac{1}{n}\sum_{i=1}^{n}\nabla f(x;\xi^{k}_{i})-\nabla f(x)\|^{2}]\leq\frac{\sigma^{2}}{n}$ also improves with $n.$ There are many other compressed methods that also improve with $n,$ including DIANA (Mishchenko et al., 2019), Accelerated DIANA (Li et al., 2020), MARINA (Gorbunov et al., 2021), DASHA (Tyurin and Richtárik, 2023a), and FRECON (Zhao et al., 2021).
@@ 1.2 Contributions
Theorem 1.6 (Informal Formulation of Theorems 4.2 and F.1).
@@
Because of the $\min$, the bound shows that it is possible to improve either the dependence on $d$ or the dependence on $\nicefrac{{\sigma^{2}}}{{\varepsilon}}$ as the number of workers $n$ increases, but not both simultaneously. The lower bound is matched either by Batch QSGD or by the non-distributed SGD method (without any communication or cooperation). Moreover, if $\tau_{\textnormal{s}}\simeq\tau_{\textnormal{w}},$ the lower bound becomes $\tilde{\Omega}\left(\min\left\{h\left(\frac{\sigma^{2}}{n\varepsilon}+1\right)\frac{L\Delta}{\varepsilon}+{\color[rgb]{0.92578125,0.578125,0.0546875}\definecolor[named]{pgfstrokecolor}{rgb}{0.92578125,0.578125,0.0546875}\tau_{\textnormal{s}}d\frac{L\Delta}{\varepsilon}},h\frac{L\Delta}{\varepsilon}+{\color[rgb]{0.92578125,0.578125,0.0546875}\definecolor[named]{pgfstrokecolor}{rgb}{0.92578125,0.578125,0.0546875}h\frac{\sigma^{2}L\Delta}{\varepsilon^{2}}}\right\}\right),$ which can be matched by Batch Synchronized SGD with the complexity (2) (without compression) or by the non-distributed SGD method. In other words, if $\tau_{\textnormal{s}}\simeq\tau_{\textnormal{w}},$ then using methods with random sparsification compression in the distributed centralized setting offers no advantage. However, if $\tau_{\textnormal{s}}\lesssim\tau_{\textnormal{w}},$ the compression techniques can help on the workers side in the regimes when $\nicefrac{{\tau_{\textnormal{w}}d}}{{n}}+\sqrt{\nicefrac{{d\tau_{\textnormal{w}}h\sigma^{2}}}{{n\varepsilon}}}$ is larger than $\color[rgb]{0.92578125,0.578125,0.0546875}\definecolor[named]{pgfstrokecolor}{rgb}{0.92578125,0.578125,0.0546875}\tau_{\textnormal{s}}d$}, due to the former scaling with $n$.
@@
Improved analysis when $\tau_{\textnormal{w}}>0.$ To obtain the complete lower bound, we extended and improved the result by Tyurin et al. (2024), which was limited for our scenario and required additional modifications to finally obtain (5) (see Sections F and 5 for details).
@@ 2 Preliminaries
For better comprehension of our new idea, we now present arguably one of the most important worst-case functions by Carmon et al. (2020), which is widely used to prove lower bounds in nonconvex optimization. It has been used by Arjevani et al. (2022; 2020a) to derive lower bounds in the stochastic setting, by Lu and De Sa (2021) in the decentralized setting, by Tyurin and Richtárik (2023b; 2024); Tyurin et al. (2024) in the asynchronous setting, by Huang et al. (2022) to show the lower iteration bound for unidirectional compressed methods, and by Li et al. (2021) in problems with a nonconvex-strongly-concave structure.
@@
$\Psi(x)=\begin{cases}0,&x\leq 1/2,\\ \exp\left(1-\frac{1}{(2x-1)^{2}}\right),&x\geq 1/2,\end{cases}\quad\textnormal{and}\quad\Phi(x)=\sqrt{e}\int_{-\infty}^{x}e^{-\frac{1}{2}t^{2}}dt.$ (7)
Notice that this function has a “chain–like” structure. If a method starts from $x^{0}=0$ and computes the gradient of $F_{T},$ then the gradient will have a non-zero value only in the first coordinate (use that $\Psi(0)=\Psi^{\prime}(0)=0$). Thus, by computing a single gradient, any “reasonable” method can “discover” at most one coordinate. At the same time, if the method wants to find an $\varepsilon$–stationary point, it should eventually discover the $T$th coordinate. These two facts imply that every “reasonable” method should compute the gradient of $F_{T}$ at least $T$ times. In the construction, Carmon et al. (2020) take $T=\Theta\left(\frac{L\Delta}{\varepsilon}\right).$ This construction is a “more technical” version of the celebrated quadratic optimization construction from (Nesterov, 2018), which has similar properties. Let us define $\textnormal{prog}(x):=\max\{i~\geq~0\,|\,x_{i}\neq 0\}\,\,(x_{0}\equiv 1),$ then the following lemma is a formalization of the described properties.
@@ 2.1 Family of distributed methods
In our lower bound, we focus on the family of methods described by Protocol 1. This protocol takes an algorithm as input and runs the standard functions of the workers and the server: the workers compute stochastic gradients locally, send compressed information, the server aggregates them asynchronously and in parallel, and sends compressed information back based on the local information. For now, we ignore the communication times from the workers to the server in Protocol 1.
@@
This is the standard assumption (Carmon et al., 2020) that generalizes the family of methods working with the span of vectors (Nesterov, 2018) and holds for the majority of methods, including GD, Adam (Kingma and Ba, 2015), DORE (Liu et al., 2020), EF21-P (Gruntkowska et al., 2023), MARINA-P, and M3 (Gruntkowska et al., 2024).
@@ 2.2 Previous Lower Bound in the Heterogeneous Setting
6:  for $k=0,1,\dots$ do
7:   Algorithm AA calculates a new point $x$ based on local information $I_{i}$: (takes 0 seconds)any vector $x\in\mathbb{R}^{d}$ such that $\textnormal{supp}(x)\in\cup_{y\in I_{i}}\textnormal{supp}(y)$ ($\textnormal{supp}(v):=\{\,i\in[d]:v_{i}\neq 0\,\}$)
[…]
9:   Add $\nabla f(x;\xi)$ to $I_{i}$ (takes 0 seconds)
10:  end for
11:end for
12:for $i=1,\dots,n$ (in parallel on the workers) do
13:  for $k=0,1,\dots$ do
14:   Algorithm AA calculates a new point $\bar{s}^{k}_{i}$ based on local information $I_{i}$: (takes 0 seconds) any vector $\bar{s}^{k}_{i}\in\mathbb{R}^{d}$ such that $\textnormal{supp}(\bar{s}^{k}_{i})\in\cup_{y\in I_{i}}\textnormal{supp}(y)$
15:   Send $\bar{\mathcal{C}}^{k}_{i}(\bar{s}^{k}_{i})$ to the server (takes $\tau_{\textnormal{w}}\times\bar{P}^{k}_{i}$ seconds, where $\bar{P}^{k}_{i}$ is the number of coordinates retained by $\bar{\mathcal{C}}^{k}_{i}(\bar{s}^{k}_{i})$)
16:   Add to $\bar{\mathcal{C}}^{k}_{i}(\bar{s}^{k}_{i})$ to II (takes 0 seconds)
17:  end for
@@
19:for $i=1,\dots,n$ (in parallel on the server) do
20:  for $k=0,1,\dots$ do
21:   Algorithm AA calculates a new point $s_{i}^{k}$ based on local information II: (takes 0 seconds)any vector $s_{i}^{k}\in\mathbb{R}^{d}$ such that $\textnormal{supp}(s_{i}^{k})\in\cup_{y\in I}\textnormal{supp}(y)$
22:   Algorithm AA compresses the point: $\mathcal{C}^{k}_{i}(s_{i}^{k})\quad\forall i\in[n]$ (takes 0 seconds)
23:   Send $\mathcal{C}^{k}_{i}(s_{i}^{k})$ to the worker $i$ (takes $\tau_{\textnormal{s}}\times P^{k}_{i}$ seconds, where $P^{k}_{i}$ is the number of coordinates retained by $\mathcal{C}^{k}_{i}(s_{i}^{k})$)
24:   Add to $\mathcal{C}^{k}_{i}(s_{i}^{k})$ to $I_{i}$ (takes 0 seconds)
25:  end for
26:end for(a vector may be added to II or $I_{i}$ at the same time as the algorithm calculates a new point; in this case, the protocol adds the vector first (with no delay since the operation takes 0 seconds))
@@
The next step in (Gruntkowska et al., 2024), in the proof of the lower bound theorem, was to analyze Protocol 1. They consider [^4] In general, they presented a more general setting where the server can zero out coordinates with any probability, capturing not only RandKK with $K=1$ and $p=\nicefrac{{1}}{{d}},$ but also RandKK with $K>1$ and other compressors. RandKK with $K=1.$ Then, since the compressor sends only one coordinate with probability $p=\nicefrac{{1}}{{d}},$ the probability that the server sends the last non-zero coordinate to the worker responsible for the current block of (6) that can progress to the new coordinate is also $p.$ Thus, the number of consecutive coordinates that the server has to send to the workers is at least $\sum_{j=1}^{T}\eta_{j},$ where $\eta_{j}$ is a geometric-like random variable with $\mathbb{P}\left(\eta_{j}=m\middle|\eta_{j-1},\dots,\eta_{1}\right)\leq p(1-p)^{m-1}$ for all $m\geq 1.$ Using classical tools from statistical analysis, one can show that $\sum_{j=1}^{T}\eta_{j}\gtrsim\nicefrac{{T}}{{p}}\simeq d\nicefrac{{L\Delta}}{{\varepsilon}}$ with high probability. Thus, under Assumption 1.4, the communication time complexity cannot be better than $\Omega\left({\color[rgb]{0.92578125,0.578125,0.0546875}\definecolor[named]{pgfstrokecolor}{rgb}{0.92578125,0.578125,0.0546875}\tau_{\textnormal{s}}d\nicefrac{{L\Delta}}{{\varepsilon}}}\right),$ which does not improve with $n.$